# FedDRL: A Trustworthy Federated Learning Model Fusion Method Based on Staged Reinforcement Learning

### 21 Jan 2024 v3

@@ 3.2 Trusted Client Selection Method
Since we use an A2C distributed reinforcement learning-based approach, as shown in Figure 1, the A2C architecture consists of a central node and K workers. Each worker contains an Actor and a Critic network, where the Actor-network generates action A, and the Critic network evaluates the action and gives the corresponding reward. Meanwhile, each worker independently interacts with the related environment to achieve sampling and training of the Actor and Critic networks. In addition, the Actor and Critic networks of the central node are used to synchronize the network information of each worker and to achieve the fusion and sharing of network parameters of multiple workers.
Figure 1: The Algorithm flow of FedDRL
@@
Therefore our main objective is to train Actor and Critic networks. We define the Actor-network parameters as $\pi(\theta)$ and the Critic network parameters as $V(w)$. Our working process for each worker and the central node is as follows. The process is shown in algorithm 1.
@@
Step 1: First, each worker initializes the local network by pulling the global network model parameters. Then, both networks train their respective Actor and Critic networks by interacting with the environment independently. Finally, the trained networks are uploaded to the central node.
@@
Step 2: After the central node collects the network parameters uploaded by all workers, it updates the global model by weighted averaging. Finally, it sends it to each worker to update the network parameters.
@@
The training process of the step 1 neutralization network is as follows. The gradient of the primary communication algorithm of the policy network is calculated as equation 8.
@@
The Loss function of the Actor-network is calculated as in equation 10, and The Critic network is calculated as in equation 11.
@@
Finally, the network parameters of the central node are calculated by the weighted average method. The process is equation 13.
@@
Algorithm 1 FedDRL Stage 1 Algorithm Process
@@
3.3 Adaptive weights calculation method
In this phase, we implement a distributed reinforcement learning approach based on TD3 [25]. In this phase, our main objective is to achieve the optimal weight assignment for each model. Since the weight values of the client models correspond to a continuous action space, we use the TD3 algorithm to accomplish this process. For each communication round in federation learning, we assume that K plausible client models are screened in the previous stage, where the parameters of the client models are set. The number of samples corresponding to the clients is defined as. These two terms are fixed in each federal communication round. We need to assign the weight values corresponding to each model to the k models based on the quality of the models, which is defined. We validate the accuracy Acc of each model based on the retained data. We need to train the reinforcement learning intelligences in each communication round of federation learning and achieve the global model fusion by assigning the optimal model weights by this intelligences. The process is described as follows.
(1) We define the t-th time, the action adopted by the agent to assign the weights of each model as equation 14.
@@ 3.3 Adaptive weights calculation method
$\theta_{global}^{k}=\sum_{\mathrm{i}=1}^{\mathrm{t}}w_{i}^{t}\theta_{i}$ (15)
(3) We use the global model aggregated in step 2 to calculate the accuracy of this model and the reward value based on the accuracy rate. And the reward value is returned to the intelligence. The intelligence is trained by steps 1-3 until the reward value of the intelligence is stable. Finally, the optimal weights of each model are output by the intelligent body. We model the process as follows.
State: We define the state of the teeth reinforcement learning as including the number of data samples S for each client, the accuracy Acc of each client’s model, and the global model Acc based on this aggregation. The process is as in 16.
@@
Action: In each stage, we need to reassign each model’s weights based on the model’s quality. The action space is shown as 17. $a_{i}^{t}$ denotes the weight value assigned to the i-th client in the state t, while the sum of the corresponding weight values of all clients is 1.
@@
Reward: We define the model accuracy aggregated using the weighted average method as $Acc_{all}$. Also, we define the global model accuracy obtained by the overall model weights assigned to the client for the mth time as $Acc_{m}$. We calculate the reward value by subtracting the difference of $Acc_{m}$ from $Acc_{all}$. If the calculated result is greater than zero, this indicates that the weights assigned by the agent improve the accuracy of the global model, and we give a positive reward. Conversely, we give a penalty reward. $\varphi$, $\phi$ denotes the reward and penalty factors, respectively. So the reward is defined as equation 18.
$Reward=\begin{cases}\varphi*(Acc_{m}-Acc_{all}),Acc_{m}>Acc_{all}\\ \phi*(Acc_{m}-Acc_{all}),Acc_{m}<Acc_{all}\end{cases}$ (18)
We use multiple parallel environments to accelerate the sampling to speed up the training agent in this phase. The stage is shown in figure 1. This stage includes a central Learner and multiple Worker nodes. Each Worker corresponds to a parallel environment. The workflow of each Worker is as follows: first, each Worker performs global model fusion based on the assigned weights; then verifies the accuracy of the global model by interacting with the parallel environment; and finally receives the reward values from the parallel environment feedback. Finally, each Worker stores the corresponding ones in the sampling buffer pool. Multiple workers interact with each environment independently, thus achieving parallel sampling to improve the sampling efficiency. After each worker collects a certain batch of samples, Learner trains the agent by taking a certain amount of sample data from the experience pool.
We use the TD3 algorithm for this phase of the method design. The TD3 algorithm consists of six network models, including an Actor-network $P(w)$, two Critic networks $Q_{1}\left(\theta_{1}\right),Q_{2}\left(\theta_{2}\right)$, and a target Actor-network $P^{\prime}(w)$, two target Critic networks $Q^{\prime}_{1}\left(\theta_{1}\right),Q^{\prime}_{2}\left(\theta_{2}\right)$. The network results of the TD3 algorithm are shown in figure 1. The learner randomly draws N batches of sample data from the buffer pool every certain round for training the model. The training processes are as follows. The process is shown in algorithm 2
@@
(3) Select the action based on the Actor-network, input the state and action into the Critic network separately, and let these two networks output the corresponding prediction reward sum.
@@
Algorithm 2 FedDRL Stage 2 Algorithm Process
@@
18:     Update Actor Network: $\nabla_{w}J\left(w\right)=N^{-1}\sum\nabla_{w}Q_{\theta_{1}}\left(s,a\right)\mid_{a=P\left({s}\right)}\nabla_{w}P\left({s}\right)$
@@
20:     Update Target Actor Network: $w^{\prime}\leftarrow\tau w+\left(1-\tau\right)w^{\prime}$
@@ 4 SYSTEM DESIGN
We propose a trusted federated model fusion framework(FedDRL) to solve the client weight assignment problem. The main workflow is divided into five steps. Figure 2 shows the system architecture of FedDRL.
@@
Step 1: First, each client downloads the global model, initializes the model parameters using the global model, and then trains the model using local private data.
Step 2: After each client’s training, each client uploads its respective model parameters to the server side.
Step 3: After receiving the client’s model parameters, the server side performs the trusted client selection and adaptive weight assignment in two stages according to the FedDRL algorithm.
In stage 1, the server uses the SelectTrustClient(.) algorithm to train the agent, then updates each network parameter of the agent, and finally outputs the optimal action to achieve the selection of trusted clients.
In stage 2, the server uses the AdaptCalculateWeight(.) algorithm to train the agent, then updates each network parameter of the agent, and finally achieves the optimal weight assignment for each client through the agent.
Step 4: The server side fuses the client models based on the calculated optimal weights and finally fuses the global models.
Step 5: The server side distributes the global model to each client.
We set several communication rounds and repeat steps 1-5 until the global model converges. The process is shown in Algorithm 3.
@@
5:     $M_{i}\leftarrow\operatorname{GetGlobalModel}(round=i)$ // Get global model and init cleint model
@@
12:     Update the $\operatorname{SelectTrustClient}\left(.\right)$ Algorithm parameters // According Algorithm 1
@@
15:     Update the $\operatorname{AdaptCalculateWeight}\left(.\right)$ Algorithm parameters // According Algorithm 2
@@ 5.1.1 Experiment datasets
We validate the FedDRL framework on three image classification datasets.
Fashion-MNIST dataset: Fashion-MNIST consists of 70,000 samples in 10 classes.
CIFAR-10 dataset: CIFAR-10 dataset consists of 60000 images in 10 classes.
CIFAR-100 dataset: CIFAR-100 has 100 classes containing 600 images for each class.
Data set partitioning: To simulate the distribution of non-IID datasets on the clients, we use the method in the paper [26] to partition different clients’ data on the open-source dataset using the Dirichlet function. The Dirichlet function could simulate the data partitioning of each client by setting the alpha parameter. When the alpha parameter tends to zero, the data distribution of clients is assigned to only a certain class of the public dataset; when the alpha parameter tends to positive infinity, the closer the distribution of client data, which is the amount and class of data is consistent across clients. To simulate the distribution of non-iid data, we set alpha to 1 to split the three datasets into ten clients. The data distribution is shown in Figure 3.
Refer to caption (a) fashionmnist
@@ 5.1.2 Comparison of methods
We compared the FedDRL algorithm with two traditional federated learning algorithms.
FedAvg[1]: The FedAvg method is the baseline of federated learning, which allocates the weight of the client model by calculating the ratio of the number of samples from each client to the total number of samples.
FedProx[2]: this algorithm improves the FedAvg method by adding a regularization term to the client model to optimize the performance of the federated learning.
@@ 5.1.3 Experimental metrics
[…]
$A_{\text{gobal }}=\left\{A_{\text{gobal }}^{(1)}\ldots A_{\text{gobal }}^{(t)}\ldots A_{\text{gobal }}^{(m)}\right\}$ (24)
To demonstrate the behavior of each client. We also calculated the accuracy of the local model for each client in each round, which is expressed as 25.
$A_{\text{c}}=\left\{A_{\text{c}}^{(1)}\ldots A_{\text{c}}^{(t)}\ldots A_{\text{c}}^{(m)}\right\},c\in\left[1,2...n\right]$ (25)
@@ 5.1.4 Experimental configuration
Hardware configuration: We completed the experiments using a high-performance workstation with an Intel i9-12900k processor, 64GB RAM, and an NVIDIA RTX3090 graphics card.
Software configuration: We used two frameworks to complete the Federated Learning and Reinforcement Learning experiments. We use the FedBolt, our self-developed experimental framework, to conduct the Federated Learning experiments. We use this framework to simulate a Federated Learning experiment environment with a different number of clients and different data distributions. For the training process of the reinforcement learning model, we use the stablebaseline3 framework [27], based on which we design two algorithms for trusted client selection and model weight assignment.
Federeted Learning Network setup: We designed different network models for different datasets. For the CIFAR-10 and CIFAR-100 datasets, we create a 6-layer CNN network to train the model. We designed a 4-layer MLP network for the FASHION-MNIST dataset.
DRL Network setup: Since we use a staged reinforcement learning strategy, we use a distributed A2C algorithm in the first stage. Each worker and the central node include a 6-layer MLP for the Actor and a 6-layer MLP for the Critic Network. In the second stage, we used the TD3 algorithm and a 6-layer MLP network for each module of TD3.
@@ 5.2 Experimental Results
We validate the FedDRL algorithm in three experiments. These include client attack experiments, low-quality model fusion experiments, and hybrid experiments. The client attack experiments focus on validating our proposed trusted client selection algorithm of stage 1. The low-quality model experiment mainly validates our proposed adaptive weight calculation algorithm of stage 2. The hybrid experiment contains both client attack and low-quality model components and mainly verifies the overall performance of the FedDRL method.
@@ 5.2.1 Malicious Client Attack Experiment
We define three attack types to simulate the client-side attack behavior in federated learning. To accomplish the experiments, we simulated different numbers of clients and different types of attacks in three datasets. The experimental configuration of the client attack is shown in table 1.
@@
Type 2: We use falsified data to perform the attack. We use a certain percentage of forged data to participate in model training (e.g., mix the cifar10 dataset with $80\%$ of cifar100 data and generate these cifar100 data labels as cifar10 corresponding label types). We conduct the attack by faking sample data to train the client’s local model, thus reducing the client model’s accuracy.
@@
Table 1: Client Attack Experimental Settings for Different Numbers of Clients
@@
Client=5 Type 1 Client1 7750 $A\leq 10\%$
@@
Client=10 Type 1 Client1,Client6 4222, 4938 $A\leq 10\%$
@@
Client=15 Type 1 Client1,Client6,Client11 3670, 3314, 4454 $A\leq 10\%$
@@
According to the experimental setup, we compared the FedDRL algorithm with the two baseline algorithms. In the attack experiments, we set the total number of communication rounds to 100 rounds, and each node performs local model training with 1 epoch. To show the attack behavior of each client and the accuracy of different algorithms in each attack scenario in more detail, we counted the accuracy of each client’s local model and the accuracy of the server-side global model in each communication round. The specific experimental results are shown in table 2.
Table 2: Global model accuracy for different number of client attack types
@@
To show the effect of the FedDRL algorithm on global model fusion at each communication round. We conducted experiments using the CIFAR10 dataset on 5, 10, and 15 clients. And We compared FedDRL with FedAvg and FedProx algorithms for the global model accuracy.
@@
We analyze the experimental results for different numbers of client models and different client data. In attack type 1, our algorithm outperforms the FedAvg and FedProx algorithms on all three data sets. In malicious data attack type 2, our algorithm outperforms the FedAvg algorithm and slightly underperforms the FedProx algorithm alone. In the attack type 3 scenario, our algorithm outperforms the comparison algorithm in most cases, especially when multiple malicious clients are involved in model fusion. To show the relationship between the global and client model’s accuracy in each attack scenario. We conducted more detailed experiments on the Cifar10 dataset. The experimental results are shown in Figure 4, Figure 5, and Figure 6.
In attack type 1, our trained agent can dynamically select trusted clients for model fusion and eliminate malicious models from participating, so our algorithm has higher reliability.
In attack type 2, our algorithm is lower than FedProx because the FedProx algorithm uses control parameters to force the models of all client nodes to converge to the global model, which will improve the accuracy of the global model by improving the accuracy of the malicious model to some extent. Our trained reinforcement learning agent will filter out low-accuracy models to participate in the fusion after several communication rounds.
In attack type 3, the trained agent can dynamically adjust the trusted clients, so our algorithm works reliably even in the case of rotating attacks by malicious clients.
Refer to caption (a) 5 clients type1 FedAvg
Refer to caption (b) 5 clients type1 FedProx
Refer to caption (c) 5 clients type1 FedDRL
Refer to caption (d) 10 clients type1 FedAvg
Refer to caption (e) 10 clients type1 FedProx
Refer to caption (f) 10 clients type1 FedDRL
Refer to caption (h) 15 clients type1 FedAvg
Refer to caption (i) 15 clients type1 FedProx
Refer to caption (j) 15 clients type1 FedDRL
@@
Figure 4: Global model accuracy in client attack type 1
Refer to caption (a) 5 clients type3 FedAvg
Refer to caption (b) 5 clients type3 FedProx
Refer to caption (c) 5 clients type3 FedDRL
Refer to caption (d) 10 clients type3 FedAvg
Refer to caption (e) 10 clients type3 FedProx
Refer to caption (f) 10 clients type3 FedDRL
Refer to caption (h) 15 clients type3 FedAvg
Refer to caption (i) 15 clients type3 FedProx
Refer to caption (j) 15 clients type3 FedDRL
@@
Figure 5: Global model accuracy in client attack type 2
Refer to caption (a) 5 clients type3 FedAvg
Refer to caption (b) 5 clients type3 FedProx
Refer to caption (c) 5 clients type3 FedDRL
Refer to caption (d) 10 clients type3 FedAvg
Refer to caption (e) 10 clients type3 FedProx
Refer to caption (f) 10 clients type3 FedDRL
Refer to caption (h) 15 clients type3 FedAvg
Refer to caption (i) 15 clients type3 FedProx
Refer to caption (j) 15 clients type3 FedDRL
@@
Figure 6: Global model accuracy in client attack type 3

### 19 Mar 2024 v4

@@ 3.2.1 Adaptive client selection method
Once we have defined the base elements of reinforcement learning, We use a distributed A2C approach to train the agent; A2C is an improved method-based A3C algorithm [26]. Figure 1 shows the A2C architecture, which consists of a central node and $K$ workers. Each worker contains an Actor and a Critic network, where the actor network generates action, and the Critic network evaluates the action and gives the corresponding reward. Meanwhile, each worker independently interacts with the related environment to achieve sampling and training of the Actor and Critic networks. In addition, the Actor and Critic networks of the central node are used to synchronize the network information of each worker and to achieve the fusion and sharing of network parameters of multiple workers.
@@
Therefore, our main objective is to train Actor and Critic networks. We define the Actor-network parameters as $\pi(\theta)$ and the Critic network parameters as $V(w)$. The process of the worker and the central node is as follows.
@@
Step 1: Each worker initializes the local network by pulling the global network model parameters from the centre node. Then, each worker trains the Actor and Critic networks by interacting with the environment independently. Finally, the two networks are uploaded to the central node.
@@
Step 2: After the central node collects the network parameters uploaded by all workers, it updates the global model by the weighted averaging method. Then, the server sends the two networks to each worker.
@@
The training process for the step 1 neutralization network is as follows: The gradient of the primary communication algorithm of the policy network is calculated as equation 4.
@@
The Loss function of the actor network is calculated as in equation 6, and The Critic network is calculated as in equation 7.
@@
Finally, each worker uploads the Actor and Critic network to the server. Then, the network parameters of the server are calculated using the weighted average method. The process is equation 9.
@@
When the parameters of the Actor and Critic networks in the central stage are updated, the central node sends down these two networks to all workers, and each worker uses the updated networks to continue interacting with the external environment. The process is repeated for the specified number of rounds until the agent at the central node can obtain a stable reward value. The process is shown in algorithm 1.
Algorithm 1 The process of trustworthy client selection
@@
3.2.2 Adaptive model weight calculation method
In this phase, our main objective is to achieve the optimal weight assignment for each model. For each communication round, we assume that $K$ trustworthy client models were selected. We need to train the agents in each communication round and use the weight output of the agent to achieve the global model fusion. We first describe the process of global model fusion for agent-based actions. We define $\theta_{i}$ as the $i$-th client model, and the all client models as $\{\theta_{1},\theta_{2},\cdots,\theta_{k}\}$. We also define $s_{i}$ as the number of samples of $i$-th client. The process is as follows:
(1) In this step, the agent needs to output the weight values for each model. We define the $t$-th time, the action adopted by the agent as equation 10.
@@ 3.2.2 Adaptive model weight calculation method
$\theta_{global}^{k}={\textstyle\sum_{t=1}^{T}}w_{i}^{t}\theta_{i}$ (11)
We aim to train the agent so that it can output the optimal fusion weight values based on the quality of each model. To accomplish this goal, we first describe the basic elements of reinforcement learning as follows:
Environment: The external environment is the server-side global model fusion module, which fuses the global model based on the actions output by the agent and then verifies the accuracy of the global model on the reserved dataset on the server side. Finally, the server side feeds back to the agent the corresponding reward and punishment values based on the accuracy of the global model.
State: We define the agent’s state information to include the number of samples corresponding to each client, the accuracy of each client’s model, and the accuracy of the global model fused using the weights output by the agent. as shown in 12.
@@
The $acc_{global}^{t}$ is the accuracy of the global model fused using the weights output by the agent.
Action: In each stage, the agent needs to assign each model’s weights based on the model’s quality. The action space is shown as 13. $a_{i}^{t}$ denotes the weight value assigned to the $i$-th client in the state t, while the sum of the corresponding weight values of all clients is 1.
@@
Reward: We define the model accuracy aggregated using the average method as $Acc_{all}$, where each model weight is $\frac{1}{N}$. At the $m$-th time, we define the weight set output by the agent as $W$. Then, we use the weight set to fusion the global model, and we define the accuracy of the global model as $Acc_{m}$. We calculate the reward value by subtracting the difference of $Acc_{m}$ from $Acc_{all}$. If the calculated result is greater than zero, this indicates that the weights assigned by the agent improve the accuracy of the global model, and we give a positive reward. Conversely, we give a penalty reward. $\varphi$, $\phi$ denotes the reward and penalty factors, respectively. So, the reward is defined as equation 14.
$Reward=\begin{cases}\varphi\cdot(Acc_{m}-Acc_{all}),Acc_{m}>Acc_{all}\\ \phi\cdot(Acc_{m}-Acc_{all}),Acc_{m}\leq Acc_{all}\end{cases}$ (14)
When we have finished defining the basic elements, We implement a distributed reinforcement learning approach based on TD3 [8] to train the agent. The training process is shown in figure 1. This stage includes a central Learner and multiple Worker nodes. Each worker corresponds to a parallel environment. The workflow of each worker is as follows: first, each worker performs global model fusion based on the assigned weights; then verifies the accuracy of the global model by interacting with the parallel environment; and finally receives the reward values from the parallel environment feedback. Finally, each worker stores the corresponding ones in the sampling buffer pool. Multiple workers interact with each environment independently, thus achieving parallel sampling to improve the sampling efficiency. After each worker collects a certain batch of samples, the Learner trains the agent by taking a certain amount of sample data from the experience pool.
The TD3 algorithm consists of six network models, including an Actor network $P(w)$, two Critic networks $Q_{1}\left(\theta_{1}\right)$,$Q_{2}\left(\theta_{2}\right)$, and a target Actor-network $P^{\prime}(w)$, two target Critic networks $Q^{\prime}_{1}\left(\theta_{1}\right),Q^{\prime}_{2}\left(\theta_{2}\right)$. Each network is shown in figure 1. The Learner randomly draws $N$ batches of sample data from the buffer pool every certain round to train the model. The training processes are as follows.
@@
(3) Select the action based on the actor network, input the state and action into the critical network separately, and let these two networks output the corresponding prediction reward sum.
@@
Repeating the above steps for the specified number of rounds, we will get the trained agent. Finally, we output the optimal value of each model through the agent. The process is shown in algorithm 2
Algorithm 2 The process of model weight calculation
@@
18:     Update Actor-Network: $\nabla_{w}J\left(w\right)=N^{-1}\sum\nabla_{w}Q_{\theta_{1}}\left(s,a\right)\mid_{a=P\left({s}\right)}\nabla_{w}P\left({s}\right)$
@@
20:     Update Target Actor-Network: $w^{\prime}\leftarrow\tau w+\left(1-\tau\right)w^{\prime}$
@@ 4 SYSTEM DESIGN
To establish a reliable federated learning process, we developed a framework for trustworthy federated learning (FedDRL). The framework employs a staged reinforcement learning approach to achieve trustworthy federated learning. In the first stage, we train agents to accomplish the selection of trustworthy clients to participate in global model fusion. Then, in the second stage, we also use the trained agent to dynamically adjust the fusion weights of each model and finally realize the optimal global model fusion. The framework workflow consists of six steps, as shown in Figure 2.
Step 1 (Local Model Training): Each client downloads the global model, initializes its parameters accordingly, and conducts model training using local private data.
Step 2 (Upload Model): After local model training, each client uploads its model parameters to the server.
Step 3 (Select Trustworthy Clients): Upon receiving client model parameters, the server employs the $SelectTrustClient(.)$ algorithm to train an agent. Subsequently, the trained agent selects trustworthy clients.
@@
Step 4 (Assigning Model Weights): The server Utilizes models from trustworthy clients and performs global model fusion. It then employs the $AdaptCalculateWeight(.)$ algorithm to train an agent, which optimizes weight assignments for each client model.
Step 5 (Fusing Global Model): The server fuses the global model using the calculated weights from the previous step.
Step 6 (Distribute Global Model): The server disseminates the global model to all clients, initiating the subsequent federation task.
The federation task is set to execute a specified number of communication rounds until the final global model is obtained. This process is shown in Algorithm 3.
@@
5:     $M_{i}\leftarrow\operatorname{GetGlobalModel}(round=i)$ // Get the global model and init client model
@@
12:     Update the $\operatorname{SelectTrustClient}\left(.\right)$ Algorithm parameters // According to Algorithm 1
@@
15:     Update the $\operatorname{AdaptCalculateWeight}\left(.\right)$ Algorithm parameters // According to Algorithm 2
@@ 5.1.1 Experiment datasets
We evaluated the FedDRL framework using three distinct image classification datasets:
Fashion-MNIST: This dataset includes 60,000 training samples and 10,000 test samples, each a 28x28 grayscale image, classified into one of 10 categories.
CIFAR-10: The CIFAR-10 dataset comprises 60,000 32x32 colour images, evenly distributed across ten classes, each containing 6,000 images.
CIFAR-100: Similar in size to CIFAR-10 but with a broader spectrum, CIFAR-100 features 100 classes with 600 images each, totalling 60,000 colour images.
Data Set Partitioning: For simulating non-IID data distribution among clients. We utilized the Dirichlet function to segregate data across various clients in the open-source dataset. This method can partition the data for each client by adjusting the alpha parameter. As the alpha parameter approaches zero, clients’ data distributions are skewed towards specific classes within the dataset. Conversely, as alpha increases towards infinity. Using the CIFAR-10 dataset as a case study, we set alpha to 1, thereby dividing the three datasets among ten clients. In the figure, Different categories are represented by distinct colours, and the length of each segment within the graphs reflects the sample count within that category. The resulting data distribution is illustrated in Figure 3.
Refer to caption (a) Fashion-MNIST
@@ 5.1.2 Comparison of Methods
We contrasted the FedDRL algorithm with two established federated learning approaches.
FedAvg[2]: Serving as the foundational benchmark in federated learning, the FedAvg method determines the weight of each client model based on the proportion of samples contributed by the client relative to the aggregate sample size.
FedProx[4]: Enhancing the FedAvg approach, FedProx incorporates a regularization term within the client model, thereby refining federated learning performance.
@@ 5.1.3 Experimental Metrics
We employed accuracy as the metric to gauge the performance of the global model in multi-classification tasks and across individual clients. Assuming n clients engage in model fusion with m communication rounds, the accuracy of the global model in the t-th round is denoted as $A_{\text{global}}^{(t)}$. The collective global model accuracies across all rounds are represented as follows:
$A_{\text{global}}=\{A_{\text{global}}^{1},A_{\text{global}}^{2},\ldots,A_{\text{global}}^{m}\}$ (20)
We denote the accuracy of the $c$-th client’s model as $A_{c}$. Additionally, we document each client model’s accuracy per round, compiling these as follows:
$A_{\text{c}}=\{A_{\text{c}}^{1},A_{\text{c}}^{2},\ldots,A_{\text{c}}^{m}\},c\in[1,2,\ldots,n]$ (21)
@@ 5.1.4 Experimental Configuration
Hardware Configuration: The experiments were conducted on a workstation equipped with an Intel i9-12900k CPU, 64GB RAM, and an NVIDIA RTX3090 GPU.
Software Configuration: We utilized two distinct frameworks for the Federated Learning and Reinforcement Learning experiments. Federated Learning trials were carried out using FedBolt, our custom-built framework, enabling simulation of varied client numbers and data distributions. For reinforcement learning model training, we employed the stablebaseline3 framework, designing two distinct algorithms for trusted client selection and model weight assignment.
Network Setup: We implemented different network architectures tailored to each dataset. For CIFAR-10 and CIFAR-100, a 6-layer CNN was utilized for model training. Conversely, a 4-layer MLP was developed for the Fashion-MNIST dataset.
Agent Network Setup: Implementing a staged reinforcement learning strategy necessitated the training of two distinct agents. The initial phase, adhering to Section 3.2.1, utilizes the A2C algorithm, with each worker and the central node comprising a 6-layer MLP actor and critic network. For the second phase, the TD3 algorithm outlined in Section 3.2.2 was employed for agent training, where each module within the TD3 setup incorporates a 6-layer MLP, with further details available in Section 3.2.2.
@@ 5.2 Experimental Results
We evaluate the FedDRL framework through four experiments: client attack scenarios, low-quality model fusion, hybrid scenarios, and multi-agent training efficiency. The client attack experiments assess the efficacy of the trustworthy client selection algorithm (stage 1). The low-quality model fusion experiment examines the adaptive weight calculation method (stage 2). The hybrid experiments, combining client attacks and low-quality model elements, validate the comprehensive performance of FedDRL. The final experiment focuses on the training efficiency of multi-agents.
@@ 5.2.1 Malicious Client Attack Experiment
In this experiment, we define three types of client-side attacks in federated learning to evaluate our FedDRL framework under adversarial conditions. The experiment spans different client numbers and attack types across three datasets, detailed in Table 1.
@@
Type 2: We use falsified data to perform the attack. We use a certain percentage of forged data to participate in model training (e.g., mix the CIFAR-10 dataset with $80\%$ of CIFAR-100 data and generate these CIFAR-100 data labels as CIFAR-10 corresponding label types). We conduct the attack by faking sample data to train the client’s local model, thus reducing the client model’s accuracy.
@@
Table 1: Experimental setup for malicious attack scenarios
@@
Clients=5 Type 1 Client1 7750 $A\leq 10\%$
@@
Clients=10 Type 1 Client1,Client6 4222, 4938 $A\leq 10\%$
@@
Clients=15 Type 1 Client1,Client6,Client11 3670, 3314, 4454 $A\leq 10\%$
@@
According to the experimental setup, we compared the FedDRL algorithm with the FedAvg and FedProx. In the attack experiments, we set the total number of communication rounds to 100 rounds, and each client performs local model training with one epoch. To show the attack behavior of each client and the accuracy of different algorithms more detail, we counted the accuracy of each client’s local model and the accuracy of the server-side global model in each communication round. The specific experimental results are shown in table 2.
Table 2: Accuracy of each algorithm under different malicious attack scenarios
@@
To show the effect of the FedDRL algorithm on global model fusion at each communication round, we conducted experiments using the CIFAR10 dataset on 5, 10, and 15 clients. We compared FedDRL with the FedAvg and FedProx algorithms for global model accuracy.
@@
We analyze the experimental results for different numbers of client models and different client data. In attack type 1, In malicious data attack type 2, our algorithm outperforms the FedAvg algorithm and slightly underperforms the FedProx algorithm alone. In the attack type 3 scenario, our algorithm outperforms the comparison algorithm in most cases, especially when multiple malicious clients are involved in model fusion.
Refer to caption (a) 10 clients FedAvg
Refer to caption (b) 10 clients FedProx
Refer to caption (c) 10 clients FedDRL
Refer to caption (d) 15 clients FedAvg
Refer to caption (e) 15 clients FedProx
Refer to caption (f) 15 clients FedDRL
@@
Figure 4: The accuracy of the global model for different number of client in attack type 1
To show the relationship between the global and client model’s accuracy in each attack scenario. We conducted more detailed experiments on the Cifar10 dataset.
In attack type 1, the global model accuracy plummets with increasing malicious clients under FedAvg and FedProx, dropping below 40% and 20% in 10 and 15 client setups, respectively. Conversely, FedDRL’s dynamic client selection maintains higher reliability. However, Our trained agent can dynamically select trusted clients for model fusion and eliminate malicious models from participating, so our algorithm has higher reliability. The experimental results are shown in Figure 4.
In attack type 2, our algorithm is better than FedAvg but lower than FedProx. The FedProx algorithm uses control parameters to force the models of each client to converge to the global model, which will improve the global model’s accuracy by improving the malicious model’s accuracy to some extent. Our trained agent will filter out low-accuracy models to participate in the fusion after several communication rounds. The experimental results are shown in Figure 5.
Refer to caption (a) 10 clients FedAvg
Refer to caption (b) 10 clients FedProx
Refer to caption (c) 10 clients FedDRL
Refer to caption (d) 15 clients FedAvg
Refer to caption (e) 15 clients FedProx
Refer to caption (f) 15 clients FedDRL
@@
Figure 5: The accuracy of the global model for different number of client in attack type 2
Refer to caption (a) 10 clients FedAvg
Refer to caption (b) 10 clients FedProx
Refer to caption (c) 10 clients FedDRL
Refer to caption (d) 15 clients FedAvg
Refer to caption (e) 15 clients FedProx
Refer to caption (f) 15 clients FedDRL
@@
Figure 6: The accuracy of the global model for different number of client in attack type 3
In attack type 3 scenarios, the FedAvg and FedProx algorithms experience significant fluctuations in global model accuracy due to alternating attack behaviors by malicious clients. Conversely, the agent within the FedDRL framework adaptively selects trusted clients, effectively excluding malicious entities from participating in model fusion, thereby enabling the FedDRL algorithm to operate with stability. The experimental results are shown in Figure 6.